OpenStudy (anonymous):

Two hockey pucks with mass 0.1 kg slide across the ice and collide. Before the collision, puck 1 is going 13 m/s to the east and puck 2 is going 18 m/s to the west. After the collision, puck 1 is going 18 m/s to the west. What is the velocity of puck 2?

OpenStudy (anonymous):

Given: m = 0.1 kg Vi1=13 m/s E Vf1=18 m/s W Vi2= 18m/s W Vf2=? Apply conservation of momentum (assuming the surface is frictionless). Total initial momentum = Total Final momentum (m1)(Vi1) + (m2)(vi2) = (m1)(vf1) + (m1)(vf2) \(\sf \large V_{f2} = \frac{(m_1)(v_{i1})+(m_2)(v_{i2})-(m_1)(v_{f1})}{(m1)}\) P.S. Take note of your sign convention

OpenStudy (anonymous):

@Data_LG2 what is m1 and m2?

OpenStudy (anonymous):

m1 is the mass of puck 1 m2 is the mass of puck 2 m1 and m2 are equal.